OpenStudy (anonymous):

What is the function for (1,1), (2,3), (3,9), (4,27)?

OpenStudy (anonymous):

This is with triangles, by the way. Not sure if that helps anyone at all.

OpenStudy (anonymous):

it looks like you are taking powers of 3

OpenStudy (anonymous):

\(1=3^0,3=3^1.9=3^2,27=3^3\)

OpenStudy (anonymous):

so maybe try \(y=3^{x-1}\)

OpenStudy (anonymous):

I have, see the problem is, I have to graph these on an online program for school, and when I graph that it doesn't go through the graphed functions. Here, I will get a pic...

OpenStudy (anonymous):

Here: http://tinypic.com/r/35d6btw/6

OpenStudy (anonymous):

This is what it is doing, and it is making like rather difficult.